If Tunisia continues to enjoy growth at a rate of 5% per annum since 2009, what will be its per capita income in 2018?


1
Expert's answer
2022-02-17T04:37:59-0500

Per capita income is a measure of how much money is earned per person in a country or region. Per capita income may be used to calculate the average per-person income for a region and to assess the population's level of living and quality of life.

To calculate a population's total per capita growth rate during a certain time period, use the following formula:

CGR = "\frac{G}{N}"


Here, CGR is the per capita growth rate. G is the change in the size of the population, expressed as a number of individuals. You find this by taking the initial population and subtracting it from the current population. Finally, N is the initial population.

The population of Tunisia in 2009 is 10,522,000 people


The forecast real GDP of 2018 is calculated as:

GDP at end date = GDP at starting date × (1 + growth rate of GDP)years

GDP at end date(2018) = GDP at starting date(2009) × (1 + growth rate of GDP(5%))years(12)

GDP at end date(2018) = 44860000000 × (1 + 0.05)10 = $ 73072212957.2

Per capita = "\frac{73072212957.2 }{10,522,000}" = 6944.71

Therefore, the projected per capita in 2018 is $ 6944.71